OpenStudy (anonymous):

log[9](27)=__log[9](3)

OpenStudy (anonymous):

is that log(9) times 3 or log(9*3)

OpenStudy (anonymous):

log(9^3)

OpenStudy (anonymous):

since \(27=3^3\) you have \[\log(27)=\log(3^3)=3\log(3)\]

OpenStudy (anonymous):

wow, i totally forgot about the 3. thanks